Write each conic sectioning standard form. Identify the conic section.
1. 25x^2+9y-225=0
2. x^2-9y^2-36=0

Answer by stanbon(75887)

Write each conic sectioning standard form. Identify the conic section.
1. 25x^2+9y-225=0
-----
25x^2 = -9y + 225
x^2 = (-9/25)y + 9
(x-0)^2 = (-9/25)(y - 25)
-------------------------
Parabola
==========================================
2. x^2-9y^2-36=0
[(x-0)^2/36] - [(y-0)^2/4] = 1
-----
Hyperbola
==================
Cheers,
Stan H.
==================